OpenStudy (adfksfnads):

Is the expression x3•x3•x3 equivalent to x3•3•3? Why or why not? Explain your reasoning.

zepdrix (zepdrix):

Third power means we're multiplying 3 x's together.\[\large\rm \color{orangered}{x^3}\cdot\color{royalblue}{x^3}\cdot\color{purple}{x^3}\qquad=\qquad \color{orangered}{x\cdot x\cdot x}\cdot\color{royalblue}{x\cdot x\cdot x}\cdot\color{purple}{x\cdot x\cdot x}\]So in total we have 9 x's multiplying together.

zepdrix (zepdrix):

Another way we could write that,\[\large\rm x^3\cdot x^3\cdot x^3=x^9\] Is this the same as \(\large\rm x^{3\cdot3\cdot3}\) ?

OpenStudy (mathmale):

Notice how zepdrix has expressed those exponents correctly. Unfortunately, x3•x3•x3 and x3•3•3 are not correct (even tho' some people can correctly interpret them). x^3: third power of x \[x^3\] another way to write "third power of x" correctly. Please answer zepdrix' question (above).

OpenStudy (adfksfnads):

@zepdrix i think that they are the same

OpenStudy (mathmale):

Think again. How could 9 = 27?